OpenStudy (anonymous):

do you agree that the equation 4-3x=2x^2 can be written in standard form in two different ways?

mathslover (mathslover):

The standard form to write quadratic equations is - \(ax^2 + bx + c =0 \) Now, you are given with equation : \(4 - 3x = 2x^2 \\ 2x^2 + 3x - 4 = 0\) Thus, it can NOT be written in the standard form in 2 different ways. Hope it helps.